What is the volume of a rock that is added to a rectangular prism with base dimensions of 15 cm by 15 cm and it raises the water level 3 cm?
Answer by Edwin McCravy(20086)

What is the volume of a rock that is added to a rectangular prism with base dimensions of 15 cm by 15 cm and it raises the water level 3 cm?

Suppose the depth of the water before dropping the rock into it is 
h cm, so the volume before dropping in the rock is:

   V = lwh = (15)(15)(h) = 225h cm³

Then the depth of the water after dropping the rock into it is 
h+3 cm, so the volume of the water and rock together after 
dropping in the rock is:

   V = lwh = (15)(15)(h+3) = 225(h+3) cm³

So the difference between those two volumes must be the rock's
volume, so we subtract to find the difference:

  Difference = 225(h+3) - 225h = 225h + 675 - 225h = 675 cm³

So the rock's volume is 675 cm³

Edwin